OpenStudy (anonymous):

A company has 160 employees. The number of part-time workers is 7/9 the number of full-time workers. How many part-time and full-time workers does the company employ?

OpenStudy (anonymous):

suppose x is the part-time workers and y is the full-time workers. x = (7/9) y the total employees is 160: x+y=160 (7/9)y+y = 160 (16/9)y=160 y=160(9/16) -> y=90 then from x+y=160, we could get x=160-90=70

OpenStudy (anonymous):

how did you get 16/9

OpenStudy (anonymous):

7/9 y + y = 7/9 y + 9/9 y = 16/9 y

OpenStudy (anonymous):

thanks could i ask you a couple more questions

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

A computer consultant charges $250 for each hour of consultation and $175 for each hour of programming for clients. In 8 hours of work, she earned $1625. How much time did she spend on each activity?

OpenStudy (anonymous):

suppose x = hours of consulation and y = hours of programming the total hours of work is 8, so x+y = 8 She earned $1625 , so: 250x + 175y = 1625 now we have 2 equations, solve those and you'll get x and y

OpenStudy (anonymous):

i got x = 3 and y=5

OpenStudy (anonymous):

so it is trial and error

OpenStudy (anonymous):

do you understand now?

OpenStudy (anonymous):

yeap

OpenStudy (anonymous):

draw the graph for the equation 2x-3y +9=0

OpenStudy (anonymous):

2x-3y=-9 i usually find the x axis intersection and y-axis intersection first. at x=0 -> y=3 , the point is (0, 3) at y=0-> x=-9/2 , the point is (-9/2 , 0) then just connect both point

OpenStudy (anonymous):

it will be looked like this

OpenStudy (anonymous):

thanks but how do you find (-9/2) on a graph

OpenStudy (anonymous):

-9/2 = -4.5 , it's between 4 and 5

OpenStudy (anonymous):

ok thanks